Expand (2+x)^(-1/2) in ascending powers of x, up to the term x^3 . Hence approximate (2/3)^(1/2) correct to 3 significant figures


1
Expert's answer
2021-09-08T10:13:44-0400

The general formula is given by

"(1+x)^n=1+nx+{n(n-1)x^2\over 2!}+{n(n-1)(n-2)x^3\over 3!}+..."

We are given

"(2+x)^{-{1\over 2}}"

Factor out "2" in "(2+x)^{-{1\over 2}}" to get

"2^{-{1\over 2}} (1+{x\over 2})^{-{1\over 2}}"

Expand "(1+{x\over 2})^{-{1\over 2}}" in ascending powers of "x" , up to the term "x^3"

"(1+x)^n=1+nx+{n(n-1)x^2\over 2!}+{n(n-1)(n-2)x^3\over 3!}+..."

"\implies (1+{x\over 2})^{-{1\over 2}}=1+(-{1\over 2})({x\over 2})+{(-{1\over 2})((-{1\over 2})-1)({x\over 2})^2\over 2!}+{(-{1\over 2})((-{1\over 2})-1)((-{1\over 2})-2)({x\over 2})^3\over 3!}+..."

"\approx 1+(-{1\over 2})({x\over 2})+{(-{1\over 2})((-{1\over 2})-1)({x\over 2})^2\over 2!}+{(-{1\over 2})((-{1\over 2})-1)((-{1\over 2})-2)({x\over 2})^3\over 3!}"

"= 1-{x\over 4}+0.09375x^2-0.0390625x^3" up to the term "x^3"

"\implies (2+x)^{-{1\over 2}} \approx 2^{-{1\over 2}}( 1-{x\over 4}+0.09375x^2-0.0390625x^3)"

"= 0.707106781( 1-{x\over 4}+0.09375x^2-0.0390625x^3)"

"= 0.707106781-0.176776695x+0.06629126x^2-0.027621358x^3"

"({2\over 3})^{1\over 2}=({3\over 2})^{-{1\over 2}}=(1+{1\over 2})^{-{1\over 2}}"

Since "(1+{x\over 2})^{-{1\over 2}} \approx 1-{x\over 4}+0.09375x^2-0.0390625x^3"

Then "(1+{1\over 2})^{-{1\over 2}} \approx 1-{1\over 4}+0.09375(1)^2-0.0390625(1)^3 \approx 0.8046875"

"\therefore ({2\over 3})^{1\over 2} \approx 0.8046875"